### v5

# MoORE: SVD-based Model MoE-ization for Conflict- and Oblivion-Resistant Multi-Task Adaptation

Abstract
Adapting large-scale foundation models in multi-task scenarios often suffers from task conflict and oblivion. To mitigate such issues, we propose a novel “model MoE-ization” strategy that leads to a conflict- and oblivion-resistant multi-task adaptation method. Given a weight matrix of a pre-trained model, our method applies SVD to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), in which each expert corresponds to the outer product of a left singular vector and the corresponding right one. We can improve the model capacity by imposing a learnable orthogonal transform on the right singular vectors. Unlike low-rank adaptation (LoRA) and its MoE-driven variants, MoORE guarantees the experts’ orthogonality and maintains the column space of the original weight matrix. These two properties make the adapted model resistant to the conflicts among the new tasks and the oblivion of its original tasks, respectively. Experiments on various datasets demonstrate that MoORE outperforms existing multi-task adaptation methods consistently, showing its superiority in terms of conflict- and oblivion-resistance. The code is available at https://github.com/DaShenZi721/MoORE.
1 Introduction
Parameter-efficient adaptation/fine-tuning xu2023parameter plays a central role in the practical deployment of large-scale pre-training models, especially in multi-task learning (MTL) scenarios zhang2021survey ; vandenhende2021multi ; crawshaw2020multi ; vithayathil2020survey ; lyu2021multi . Take large language models (LLMs) zhao2023survey in natural language processing (NLP) tasks as an example. To increase intrinsic knowledge and maintain generalization power, a pre-trained LLM often needs to learn multiple downstream tasks in different domains simultaneously or sequentially. To achieve multi-task adaptation, parameter-efficient fine-tuning (PEFT) methods like low-rank adaptation (LoRA) hu2021lora and its variants wang2023multilora ; agiza2024mtlora ; tian2024hydralora ; wang2024multimodal ; liaohmora have been proposed. However, the practical applications of these methods are often limited because they often suffer from task conflict and oblivion: $i$) The model adapted for one task often leads to the performance degradation in other tasks (also known as negative transfer wang2019characterizing ; zhang2022survey or destructive interference crawshaw2020multi ). ii)ii) Learning new tasks may result in catastrophic forgetting vithayathil2020survey — the model performance deteriorates severely in previously learned tasks.
(a) An illustration of our SVD-based model MoE-ization strategy and MoORE architecture
(b) Resistance to task conflict
(c) Resistance to task oblivion
(d) Inference efficiency
Figure 1: (a) An illustration of our model MoE-ization strategy and the corresponding MoORE architecture. (b) The comparison for various multi-task adaptation methods on fine-tuning LLaMA-3.1 8B grattafiori2024llama3herdmodels on the CSR-MTL constructed by nine tasks allenai:arc ; clark2019boolq ; OpenBookQA2018 ; Bisk2020 ; sap2019socialiqa ; zellers2019hellaswag ; sakaguchi2021winogrande ; talmor-etal-2019-commonsenseqa . MoORE consistently works better than the baselines when the number of tasks is larger than one. (c) Before adaptation, LLaMA-3.1 8B achieves encouraging overall performance (i.e., the gray dashed line) in seven tasks hendryckstest2021 ; hendrycks2021ethics ; zhou2023instructionfollowingevaluationlargelanguage ; suzgun2022challenging ; rein2024gpqa ; chen2021evaluating ; austin2021program ; cobbe2021gsm8k . MoORE mitigates the performance degradation and outperforms the most competitive baseline LoRAMoE dou2024loramoe . (d) MoORE’s runtime is comparable to that of its competitors. Compared to the original LLaMA-3.1 8B (i.e., the gray dashed line), MoORE increases the inference time moderately.
Essentially, task conflict and oblivion arise because the diversity of different tasks requires the models to adapt their parameters in different directions yadav2023ties ; yu2024language ; deep2024della ; huang2024emr ; du2024parameter . Some recent methods liaohmora ; li2024mixlora ; tian2024hydralora combine the Mixture-of-Experts (MoE) architecture jacobs1991adaptive ; shazeer2017outrageously ; lepikhin2020gshard with PEFT, mitigating the interferences across different tasks by activating task-specific parameters. Given a layer of a pre-trained model, these methods learn multiple adapters associated with a router in multi-task scenarios. Each adapter may inherit task-specific knowledge, and the router selects/fuses these adapters based on the input data and tasks.
In principle, we call the above strategy “Model MoE-ization” because it converts the original linear layer to an MoE architecture. In this study, we propose a novel model MoE-ization strategy, with the help of the singular value decomposition (SVD), leading to a conflict- and oblivion-resistant multi-task adaptation method. As illustrated in Figure 1(a), given a weight matrix of a pre-trained model, our method applies SVD abdi2007singular to it and introduces a learnable router to adjust its singular values based on tasks and samples. Accordingly, the weight matrix becomes a Mixture of Orthogonal Rank-one Experts (MoORE), where each expert is constructed by the outer product of a left singular vector and the corresponding right one. To further improve the capacity of MoORE, we impose a learnable orthogonal adapter, which is implemented by a Householder reflection adaptation module yuan2024bridging .
Unlike existing methods, which learn some additional low-rank experts without any constraints on their relations, MoORE extracts many rank-one experts intrinsically from the SVD of the original weight matrix. Such a design imposes the orthogonality of the experts, avoiding their information redundancy and undesired interferences. Moreover, similar to existing orthogonal fine-tuning strategies qiu2023controlling ; yuan2024bridging , MoORE maintains the column space of the original weight matrix and thus inherit the pre-training ability better. Thanks to the above two properties, MoORE consistently outperforms SOTA methods in various multi-task adaptation scenarios. The results in Figure 1 show the superiority of MoORE in mitigating task conflict and oblivion and its competitive inference efficiency.
2 Related Work and Preliminaries
2.1 Adapter-based Methods for Multi-Task Adaptation
Multi-task adaptation aims to fine-tune a pre-trained foundation model simultaneously or sequentially in multiple downstream tasks liu2024mftcoder ; lin2024pemt . Focusing on this problem, many adapter-based methods have been proposed liu2022polyhistor ; raychaudhuri2022controllable ; sung2022vl ; wang2024multimodal ; lu2024prompt . In particular, denote the data of $K$ tasks as $\{\mathcal{D}_{k}\}_{k=1}^{K}$. Given a pre-trained foundation model, whose parameters are denoted as $\bm{\theta}$, multi-task adaptation can often be formulated in the framework of maximum likelihood estimation:
$\sideset{}{{}_{\Delta\bm{\theta}}}{\max}\sideset{}{{}_{k=1}^{K}}{\sum}P_{\bm{\theta}\cup\bm{\Delta\theta}}(\mathcal{D}_{k}),$ (1)
where $P_{\bm{\theta}\cup\bm{\Delta\theta}}(\mathcal{D}_{k})$ denotes the likelihood of the $k$-th task’s data, which is parameterized by the original model parameters $\bm{\theta}$ and the adapters’ parameters $\Delta\bm{\theta}$.
Unlike learning and ensembling different models, adapter-based methods improve the overall model performance by sharing parameters and domain knowledge across various tasks, leading to moderate increases in parameters and complexity. For instance, Hyperformer mahabadi2021parameter utilizes a shared hypernetwork to generate task-specific adapters, reducing the number of learnable parameters. Recently, some methods extend LoRA hu2021lora to multi-task adaptation, e.g., MultiLoRA wang2023multilora , MTL-LoRA yang2025mtl , HydraLoRA tian2024hydralora , and so on, which learns multiple low-rank adapters to handle diverse tasks.
2.2 Connections to MoE Architectures
The Mixture-of-Experts (MoE) architecture was initially introduced by the work in jacobs1991adaptive , which is constructed by multiple specialized networks (called experts) and a router. Given an input data $\bm{x}\in\mathcal{X}$, the MoE derives the output $\bm{y}\in\mathcal{Y}$ as $\bm{y}=\sum_{m=1}^{M}g_{m}(\bm{x})f_{m}(\bm{x})$, where $f_{m}:\mathcal{X}\mapsto\mathcal{Y}$ denotes the $m$-th expert, which achieves a mapping from the sample space $\mathcal{X}$ to the output space $\mathcal{Y}$. $g:\mathcal{\mathcal{X}}\mapsto\mathbb{R}^{M}$ denotes the router, and $g_{m}$ denotes its $m$-th output. The router adjusts the experts’ significance based on input data. When applying a sparse routing strategy, i.e., activating only a few experts for each input shazeer2017outrageously ; lepikhin2020gshard ; fedus2022switch ; artetxe2021efficient ; du2022glam ; mustafa2022multimodal , the MoE architecture supports building large-scale models while maintaining computational efficiency. Due to its advantages, many large language models, e.g., DeepSeek dai2024deepseekmoe , Grok3, and Qwen3111https://github.com/QwenLM/Qwen3, are built based on MoE, and many efforts have been made to convert well-trained dense models into MoE architectures komatsuzaki2022sparse ; he2024upcycling ; riquelme2021scaling ; nakamura2025drop ; zhang2024bam .
As mentioned before, most existing adapter-based multi-task adaptation methods actually “MoE-ize” pre-trained models. Given a weight matrix $\bm{W}$ and its input $\bm{x}$, these methods apply multiple low-rank adapters as experts zadouri2023pushing ; dou2024loramoe ; liu2024moe ; luo2024moelora and add them to the pre-trained models, i.e.,
$\bm{y}=\bm{Wx}+\sideset{}{{}_{m=1}^{M}}{\sum}g_{m}(\bm{x})\bm{B}_{m}\bm{A}_{m}\bm{x},$ (2)
where $\bm{A}_{m}$ and $\bm{B}_{m}$ are low-rank matrices constructing the $m$-th expert $f_{m}$. For the router $g(\bm{x})$, some attempts have been made to develop advanced routing strategies, e.g., the dynamic routing in AdaMoLE liu2024adamole and the token-task hybrid routing in HMoRA liaohmora . Recently, some new MoE architectures have been proposed, including the asymmetric “Hydra” structure in HydraLoRA tian2024hydralora , the LEMoE wang2024lemoe for lifelong model editing, and the OMoE feng2025omoe for orthogonal output. However, the above methods rely purely on data-driven strategies to determine the experts’ functionality and domain knowledge. Without necessary regularization on the relations across different experts, the experts often suffer from the load imbalance issue, i.e., a limited number of experts are over-trained and applied for multiple tasks, while many experts are seldom used. This issue harms the capacity and generalizability of the models, increasing the risk of task conflict and oblivion.
3 Proposed Method
In this study, we propose a new model MoE-ization method for multi-task adaptation. In principle, our method imposes orthogonality on the experts and further regularizes their output spaces, which helps mitigate task conflict and oblivion.
3.1 SVD-based Model MoE-ization
Consider a pre-trained weight matrix $\bm{W}\in\mathbb{R}^{D_{\text{out}}\times D}$. Without loss of generality, we assume $D_{\text{out}}\geq D$ and $\text{Rank}(\bm{W})=D$. The SVD of the matrix is denoted as
$\bm{W}=\bm{U}\text{diag}(\bm{\sigma})\bm{V}^{\top}=\sideset{}{{}_{d=1}^{D}}{\sum}\underbrace{\sigma_{d}}_{\text{weight}}\cdot\underbrace{(\bm{u}_{d}\bm{v}_{d}^{\top})}_{\text{expert}},$ (3)
where $\bm{U}=[\bm{u}_{1},\cdots,\bm{u}_{D}]\in\mathbb{R}^{D_{\text{out}}\times D}$ contains left singular vectors, $\bm{V}=[\bm{v}_{1},\cdots,\bm{v}_{D}]\in\mathbb{R}^{D\times D}$ contains right singular vectors, and $\bm{\sigma}=[\sigma_{1},\cdots,\sigma_{D}]^{\top}$ is a vector of singular values.
As shown in (3), there exists an intrinsic but static MoE architecture hidden in the SVD of the weight matrix — each $\bm{W}$ corresponds to the mixture of $D$ orthogonal and rank-one experts, in which the $d$-th expert is the outer product of $\bm{u}_{d}$ and $\bm{v}_{d}$ and its weight is fixed as $\sigma_{d}$. Motivated by this intrinsic MoE, our method derives the proposed MoORE model for multi-task adaptation, which reuses the experts while introducing the following two modifications:
  • A hybrid routing strategy: Inspired by HMoRA liaohmora , we adjust the experts’ weights according to input data and tasks, leading to a hybrid routing strategy. Given an input of the $k$-th task, denoted as $\bm{x}^{(k)}\in\mathbb{R}^{d}$, we determines the weight of the $d$-th expert as
    $g_{d}(\bm{x}^{(k)})=\underbrace{\bm{p}_{d}^{\top}\bm{t}_{k}}_{\text{task-level}}+\underbrace{\bm{q}_{d}^{\top}\bm{\Gamma}\bm{x}^{(k)}}_{\text{sample-level}},$ (4)
    where $\bm{t}_{k}\in\mathbb{R}^{D_{t}}$ denotes the embedding of the $k$-th task, and the vector $\bm{p}_{d}\in\mathbb{R}^{D_{t}}$ projects the task embedding to a task-level weight. Similarly, applying the matrix $\bm{\Gamma}\in\mathbb{R}^{D_{s}\times D}$ and the vector $\bm{q}_{d}\in\mathbb{R}^{D_{s}}$, we project the input $\bm{x}^{(k)}$ to a sample-level weight. In practice, we set $D_{s},D_{t}\ll D$ to reduce the router’s parameters and computational cost. The final weight $g_{d}(\bm{x}^{(k)})$ is the summation of the task- and sample-level weights.
  • An orthogonal adapter of input: To further increase the model capacity, we can apply a learnable orthogonal transform to the input, i.e., $\bm{H}\bm{x}$, where the learnable orthogonal transform $\bm{H}$ can be implemented efficiently by the butterfly orthogonal fine-tuning (BOFT) module in liu2024parameter , the Givens rotation adapter ma2024parameter , or the Householder reflection adapter (HRA) yuan2024bridging . In this study, we implement $\bm{R}$ by HRA, which corresponds to the product of $L$ Householder reflections, i.e.,
    $\bm{H}=\sideset{}{{}_{\ell=1}^{L}}{\prod}\Big(\bm{I}-\frac{1}{\|\bm{r}_{\ell}\|_{2}^{2}}\bm{r}_{\ell}\bm{r}_{\ell}^{\top}\Big),$ (5)
    whose learnable parameters are $\bm{R}=[\bm{r}_{1},\cdots,\bm{r}_{L}]\in\mathbb{R}^{D\times L}$. Note that applying orthogonal adapters maintains the angles between neurons (i.e., the rows of the weight matrix $\bm{W}$), which helps preserve the knowledge of the pre-trained model when enhancing model capacity qiu2023controlling ; liu2024parameter ; yuan2024bridging .
Applying the above SVD-based model MoE-ization strategy, we obtain the proposed MoORE model, which introduces $D$ orthogonal and rank-one experts into the model and encodes the input data as
$\begin{aligned} &\bm{y}=\bm{WHx}^{(k)}+\sideset{}{{}_{d=1}^{D}}{\sum}\underbrace{g_{d}(\bm{x}^{(k)})}_{\text{router}}\underbrace{(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})}_{\text{expert}}\bm{x}^{(k)}=\bm{U}\text{diag}(g(\bm{x}^{(k)})+\bm{\sigma})\bm{V}^{\top}\bm{H}\bm{x}^{(k)},\\ &\text{where}~~g(\bm{x}^{(k)})=\bm{P}^{\top}\bm{t}_{k}+\bm{Q}^{\top}\bm{\Gamma}\bm{x}^{(k)}\in\mathbb{R}^{D}.\end{aligned}$ (6)
The learnable parameters of MoORE include $\bm{T}=[\bm{t}_{k}]\in\mathbb{R}^{D_{t}\times K}$, $\bm{P}=[\bm{p}_{d}]\in\mathbb{R}^{D_{t}\times D}$, $\bm{Q}=[\bm{q}_{d}]\in\mathbb{R}^{D_{s}\times D}$, $\bm{\Gamma}\in\mathbb{R}^{D_{s}\times D}$, and $\bm{R}\in\mathbb{R}^{D\times L}$. As shown in (6), given the SVD of $\bm{W}$ (which can be computed in advance), we can implement MoORE with low complexity via the following steps:
$1)~\bm{z}=\underbrace{\bm{VHx}^{(k)}}_{\mathcal{O}(D(L+D))},\quad 2)~g(\bm{x}^{(k)})=\underbrace{\bm{P}^{\top}\bm{t}_{k}}_{\mathcal{O}(DD_{t})}+\underbrace{\bm{Q}^{\top}\bm{\Gamma}\bm{x}^{(k)}}_{\mathcal{O}(DD_{s})},\quad 3)~\bm{y}=\underbrace{\bm{U}(g(\bm{x}^{(k)})+\bm{\sigma})\bm{z}}_{\mathcal{O}(DD_{\text{out}})}.$ (7)
As shown in (7), the overall complexity of MoORE is $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}+L))$. In practice, we set $L,D_{t},D_{s}\ll\min\{D,D_{\text{out}}\}$ to reduce the complexity. Moreover, we can merge the orthogonal adapter into each expert in the inference phase, i.e., obtaining $\bm{V}^{\prime}=\bm{V}^{\top}\bm{H}$, and the complexity further reduces to $\mathcal{O}(D(D_{\text{out}}+D+D_{t}+D_{s}))$. Figure 1(d) shows that the inference efficiency of MoORE is comparable to its competitors.
Table 1: Comparisons for various MoE-based multi-task adaptation methods on their implementations and properties, where $M$ ($D$ in our method) is the number of experts, $k$ is the task index, $\text{GS}(\cdot)$ denotes the Gram-Schmidt orthogonalization bjorck1994numerics .
Method Router Experts #Experts Rank Orthogonality Maintaining
of Experts of Experts $\text{Range}(\bm{W})$
LoRAMoE dou2024loramoe $\text{Softmax}(\bm{Sx})$ $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
MixLoRA li2024mixlora $\text{Top}_{2}(\text{Softmax}(\bm{Sx}))$ $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
MoSLD zhao2025mosld $\text{Softmax}(\bm{Sx})$ $\{\bm{B}\bm{A}_{m}\}_{m=1}^{M}$ $M$ $r$ No No
HydraLoRA tian2024hydralora $\text{Softmax}(\bm{Sx})$ $\{\bm{B}_{m}\bm{A}\}_{m=1}^{M}$ $M$ $r$ No No
MTL-LoRA yang2025mtl $\text{Softmax}(\bm{\phi}_{k})$ $\{\bm{B}_{m}\bm{\Lambda}_{k}\bm{A}\}_{m=1,k=1}^{M,K}$ MKMK $r$ No No
OMoE feng2025omoe $\text{Softmax}(\bm{Sx})$ $\text{GS}(\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M})$ $M$ $r$ Yes No
MoORE (Ours) $\bm{P}^{\top}\bm{t}_{k}+\bm{Q}^{\top}\bm{\Gamma}\bm{x}$ $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$ $D$ 11 Yes Yes
3.2 Comparisons with Existing MoE-based Multi-Task Adaptation Methods
Our SVD-based model MoE-ization method provides a new technical route for multi-task adaptation: Instead of learning an MoE with few strong extrinsic experts, our method constructs an MoE with many simple but structured experts intrinsically based on the pre-trained weight matrix. Tables 1 and 2 compare different methods on their MoE architectures, theoretical properties, and computational efficiency, respectively.
  • The design of router. Given an input $\bm{x}\in\mathbb{R}^{D}$, most existing methods leverage a sample-level routing strategy. Typically, they apply a linear projection $\bm{S}\in\mathbb{R}^{M\times D}$ to it and pass the projection result through a softmax operator, leading to a nonnegative and normalized weight vector for $M$ experts. Among them, MixLoRA li2024mixlora further applies a sparse routing mechanism — for each input, it only activates the two experts that correspond to the top-2 weights. Instead, MTL-LoRA yang2025mtl leverages a task-level routing strategy, which determines the experts’ weights by passing a task-specific embedding (i.e., $\bm{\phi}_{k}\in\mathbb{R}^{M}$, $k\in\{1,\cdots,K\}$ indicates the task index) through a softmax operator. Unlike existing methods, our MoORE considers the task- and sample-level information jointly. The advantage of this hybrid routing strategy is that when the same sample serves for different tasks, MoORE can assign task-specific weights to the experts and thus leverage different domain knowledge accordingly.
  • The design of experts. Most existing methods apply $M$ low-rank adapters as experts, i.e., $\{\bm{B}_{m}\bm{A}_{m}\}_{m=1}^{M}$, where $\bm{B}\in\mathbb{R}^{D_{\text{out}}\times r}$ and $\bm{A}\in\mathbb{R}^{r\times D}$ are two rank-$r$ matrices. To reduce the number of learnable parameters, some methods, e.g., MoSLD zhao2025mosld , HydraLoRA tian2024hydralora , and MTL-LoRA yang2025mtl , reuse the same $\bm{A}$ or $\bm{B}$ for all experts. MTL-LoRA yang2025mtl further introduces a task-specific matrix $\bm{\Lambda}_{k}$ for each expert, where k=1,,Kk=1,...,K indicates the task index. As a result, it creates $MK$ low-rank experts and activates $M$ experts per task. Our MoORE contains $D$ orthogonal rank-one experts $\{\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H}\}_{d=1}^{D}$, in which the orthogonal adapter $\bm{H}$ is shared by all experts. Unlike existing methods, MoORE applies many simple but structured experts. Such a design has several advantages:
    1. 1.
      Imposing orthogonality for mitigating task conflict: The experts of MoORE are orthogonal to each other because $\bm{u}_{d}^{\top}\bm{u}_{d^{\prime}}=0$ for all $d\neq d^{\prime}$. The orthogonality ensures that the experts have different functionalities and domain knowledge, without redundant information. In particular, by activating different experts for different tasks, MoORE suppresses the interferences across the tasks in the training phase, thus mitigating the task conflict issue. It is worth noting that O-LoRA wang2023orthogonal also introduces orthogonality constraints within its LoRA adapters. However, there are two notable differences between O-LoRA and MoORE: $i)$ O-LoRA uses a regularization term to pursue orthogonality, which can not strictly make its LoRA adapters orthogonal to each other. ii)ii) O-LoRA does not adopt a MoE architecture. To deal with different tasks, O-LoRA adds new LoRA adapters for each sequentially incoming task, rather than designing and learning a flexible routing mechanism for given and fixed experts. As a result, the model complexity of O-LoRA is linear to the number of tasks. Among existing MoE-based methods, OMoE feng2025omoe is the only one imposing orthogonality on experts. However, it applies Gram-Schmidt orthogonalization algorithm bjorck1994numerics to the concatenation of $M$ experts’ output, i.e., $\text{GS}([\bm{B}_{1}\bm{A}_{m}\bm{x},\cdots,\bm{B}_{M}\bm{A}_{m}\bm{x}])$. As a result, imposing orthogonality requires additional $\mathcal{O}(D_{\text{out}}M^{2})$ operations per sample, which is less efficient than ours.
    2. 2.
      Maintaining $\text{Range}(\bm{W})$ for mitigating task oblivion: The column space of each expert, i.e., $\text{Range}(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})$, is the same with $\text{Range}(\bm{u}_{d})$. Accordingly, the output space of MoORE is the direct sum of $\{\text{Range}(\bm{u}_{d})\}_{d=1}^{D}$, which is the same as $\bm{W}$’s column space, i.e.,
      $\sideset{}{{}_{d=1}^{D}}{\bigoplus}\text{Range}(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})=\sideset{}{{}_{d=1}^{D}}{\bigoplus}\text{Range}(\bm{u}_{d})=\text{Range}(\bm{U})=\text{Range}(\bm{W}).$ (8)
      In single task adaptation scenarios, the work in yuan2024bridging has shown that maintaining the column space of the weight matrix makes the adapted model inherit the ability of the pre-trained model better, mitigating the oblivion of the previously pre-trained task. In our work, we find that such maintenance is helpful in multi-task adaptation scenarios as well.
  • Computational efficiency. Table 2 shows each method’s learnable parameters and computational complexity. For the methods using the sample-based routing strategy, their routers contain $MD$ learnable parameters. Given a sample, the complexity of the router is $\mathcal{O}(MD)$. For the method yang2025mtl using the task-based routing strategy, its router is lightweight, containing $MK$ learnable parameters and determining the weights of its experts with complexity $\mathcal{O}(M)$. For the experts, using $M$ rank-$r$ experts leads to the complexity $\mathcal{O}(M(D_{\text{out}}+D)r)$. Reusing $\bm{A}$ or $\bm{B}$ (e.g., MoSLD zhao2025mosld and HydraLoRA tian2024hydralora ) and applying sparse routing (e.g., MixLoRA li2024mixlora ) can reduce the computational complexity significantly. In contrast, introducing additional parameters (e.g., the $\bm{\Lambda}_{k}$ in MTL-LoRA yang2025mtl ) or operations (e.g., the Gram-Schmidt orthogonalization in OMoE feng2025omoe ) leads to higher complexity. Existing methods construct an MoE with $M$ rank-$r$ experts, while our MoORE is an MoE with $D$ rank-one experts, whose number of experts is determined by the input dimension and thus much larger than $M$. To improve the computational efficiency of MoORE, we set $D_{s}$ and $D_{t}$ comparable to the rank $r$ and set $L$ comparable to $M$, respectively. As a result, the number of learnable parameters and the complexity of MoORE become comparable to most existing methods.
Table 2: Comparisons for various MoE-based multi-task adaptation methods on their computational efficiency.
Method #Learnable Parameters Computational Complexity
Router Experts
LoRAMoE dou2024loramoe MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+M(D+D_{\text{out}}r+Dr))$
MixLoRA li2024mixlora MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+MD+(D_{\text{out}}r+Dr))$
MoSLD zhao2025mosld MDMD $(D_{\text{out}}+MD)r$ $\mathcal{O}(D_{\text{out}}D+D_{\text{out}}r+M(Dr+D))$
HydraLoRA tian2024hydralora MDMD $(MD_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+M(D_{\text{out}}r+D))$
MTL-LoRA yang2025mtl MKMK $(MD_{\text{out}}+Kr+D)r$ $\mathcal{O}(D_{\text{out}}D+Dr+MD_{\text{out}}r+r^{2})$
OMoE feng2025omoe MDMD $M(D_{\text{out}}+D)r$ $\mathcal{O}(D_{\text{out}}(D+M^{2})+M(D+D_{\text{out}}r+Dr))$
MoORE (Ours) $D_{t}K+(D_{t}+2D_{s})D$ LDLD $\mathcal{O}((D_{\text{out}}+D+D_{s}+D_{t}+L)D)$
3.3 Comparisons with Existing SVD-based Fine-Tuning Methods
Table 3: Comparisons for various SVD-based fine-tuning methods on their implementations.
Method Model in Training Learnable Parameters Model in Inference
SVF sun2022singular $\bm{U}\text{diag}(\sigma)\bm{V}^{\top}$ Nonparametric $\sigma$ Original architecture: $\bm{W}_{\text{new}}$
SVDiff han2023svdiff $\bm{U}\text{diag}(\text{ReLU}(\sigma+\delta))\bm{V}^{\top}$ Nonparametric $\delta$ Original architecture: $\bm{W}_{\text{new}}$
SVFT lingam2024svft $\bm{U}(\text{diag}(\sigma)+\bm{M})\bm{V}^{\top}$ Nonparametric $\bm{M}$ Original architecture: $\bm{W}_{\text{new}}$
KaSA wang2024kasa $\bm{W}_{lr}+\Delta\bm{U}\text{diag}(\Delta\sigma)\Delta\bm{V}^{\top}$ Nonparametric $\Delta\bm{U},\Delta\sigma,\Delta V^{\top}$ Original architecture: $\bm{W}_{\text{new}}$
SVFCL wang2025singular $\bm{U}\text{diag}(\sigma+\sum_{i=0}^{t}\sigma_{i})\bm{V}^{\top}$ Nonparametric $\sigma_{i}$’s Original architecture: $\bm{W}_{\text{new}}$
MoORE (Ours) $\bm{WH}$ $\sum_{d=1}^{D}g_{d}(\bm{x}^{(k)})(\bm{u}_{d}\bm{v}_{d}^{\top}\bm{H})$ Parametric $g(\cdot)$ and Householder reflections $\bm{H}$ An MoE, where $\bm{WH}$ and
$\bm{v}_{d}^{\top}\bm{H}$ are pre-computed
Table 4: Comparisons for various SVD-based fine-tuning methods on their applications.
Method Multi-Task Learning Consider Task Conflict Consider Task Oblivion
SVF sun2022singular No No No
SVDiff han2023svdiff No No No
SVFT lingam2024svft No No No
KaSA wang2024kasa No No No
SVFCL wang2025singular Yes (incremental learning) No Partially discussed
MoORE (Ours) Yes Yes, by orthogonal experts Yes, by maintaining $\text{Range}(\bm{W})$
Our method employs SVD to perform model MoE-ization, constructing a complete MoE architecture. This fundamentally differs from existing SVD-based fine-tuning methods peng2024sam ; sun2022singular ; han2023svdiff ; lingam2024svft ; wang2024kasa ; wang2025singular . To more clearly illustrate the differences between MoORE and these methods, we select five recent ones for comparison, including SVF sun2022singular , SVDiff han2023svdiff , SVFT lingam2024svft , KaSA wang2024kasa , and SVFCL wang2025singular . Tables 3 and 4 compare different methods in terms of implementation and application, respectively.
  • In the aspect of architecture, MoORE is the only method that learns a parametric router to adjust singular values, leading to an MoE architecture for inference. In MoORE, each expert is constructed by the outer product of singular vectors and the learnable Householder reflections.
  • In the aspect of application, only MoORE and SVFCL consider multi-task adaptation. MoORE is applicable for both simultaneous and sequential multi-task adaptation, while SVFCL only considers incremental learning. Moreover, MoORE makes the first attempt to connect task conflict and oblivion to the orthogonality of experts and the column space of the parameter matrix.
4 Experiments
To demonstrate the effectiveness of MoORE in multi-task adaptation, we apply three MTL datasets and conduct comprehensive experiments on them. Representative results are shown in Figure 1 and the following content. More experimental details, e.g., the basic information of datasets, hyperparameter settings, ablation studies, routing weight analysis, and numerical results associated with figures, are shown in the Appendix.
4.1 Implementation Details
Base model and baselines. In the following experiments, we utilize LLaMA-3.1 8B [^2] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct grattafiori2024llama3herdmodels as the base model and adapt it by various multi-task adaptation methods. In particular, we compare MoORE with LoRA hu2021lora and the methods incorporating low-rank adapters as MoEs, including LoRAMoE dou2024loramoe , MoSLD zhao2025mosld , MTL-LoRA yang2025mtl , HydraLoRA tian2024hydralora , and MixLoRA li2024mixlora . We implement the MoE architectures of the baselines mainly based on their default settings. For a fair comparison, we modify some baselines’ hyperparameters to make the number of learnable parameters comparable for each method. For MoORE, we MoE-ize all linear layers of LLaMA-3.1 8B, including the “QKVOQKVO” modules of attention layers and the weight matrices of FFNs.
Two datasets for evaluating conflict-resistance. We consider two MTL datasets for commonsense reasoning (CSR) and natural language understanding (NLU), respectively. The CSR-MTL dataset is constructed by nine tasks, including ARC-Challenge (ARC-C), ARC-Easy (ARC-E) allenai:arc , OpenBookQA (OBQA) OpenBookQA2018 , PIQA Bisk2020 , SocialIQA (SIQA) sap2019socialiqa , BoolQ clark2019boolq , Hellaswag (HellaS) zellers2019hellaswag , Winogrande (WinoG) sakaguchi2021winogrande , and CommonsenseQA (CSQA) talmor-etal-2019-commonsenseqa . These tasks are widely used to evaluate LLMs on various commonsense reasoning challenges, ranging from genuine grade-school level science questions to physical commonsense reasoning. The NLU-MTL dataset consists of seven tasks from GLUE wang2019glue , including CoLA, SST-2, MRPC, QQP, MNLI, QNLI, and RTE. These tasks are applied to evaluate the natural language understanding capabilities of LLMs, including natural language inference, textual entailment, sentiment analysis, semantic similarity, and so on.
One dataset for evaluating oblivion-resistance. In addition, we construct one more dataset, called OR-MTL, for evaluating the oblivion-resistance of different methods. The dataset includes seven tasks, including MMLU hendryckstest2021 ; hendrycks2021ethics , IFEval zhou2023instructionfollowingevaluationlargelanguage , BIG-Bench Hard (BBH) suzgun2022challenging , GPQA rein2024gpqa , HumanEval chen2021evaluating , MBPP austin2021program , and GSM-8K cobbe2021gsm8k . The base model, LLaMA-3.1 8B, achieves encouraging performance in these tasks. After adapting it on CSR-MTL, we record the performance of the adapted models in OR-MTL and assess the ability of different methods to mitigate task oblivion.
Experimental settings. When adapting the pre-trained model on CSR-MTL and NLU-MTL, we set the training epoch to $2$ and $5$, respectively. The learning rate is set to $3\times 10^{-4}$, with AdamW loshchilov2017decoupled as the optimizer. For CSR-MTL, we set the batch size to $8$, whereas for NLU-MTL, we set the batch size to $64$. Both training and testing are conducted on one NVIDIA A100 GPU.
Table 5: Results (%) of various methods on CSR-MTL. The best results on each dataset are shown in bold, and the second best results are shown in underline.
Method #Params ARC-C ARC-E BoolQ OBQA PIQA SIQA HellaS WinoG CSQA Overall
Full Fine-tuning 100.00% 80.12 87.92 74.56 87.60 88.96 80.55 95.91 81.69 81.16 84.27
LoRA hu2021lora 2.09% 77.56 85.77 70.43 81.60 82.97 76.00 93.00 71.11 77.40 79.54
MixLoRA li2024mixlora 3.00% 79.18 87.50 72.02 86.60 87.38 78.86 93.65 77.35 80.43 82.55
MoSLD zhao2025mosld 1.49% 80.29 86.87 74.16 86.40 88.58 79.73 95.03 80.58 80.34 83.55
HydraLoRA tian2024hydralora 2.72% 79.27 88.09 74.34 85.60 88.96 79.84 95.36 83.03 80.10 83.84
MTL-LoRA yang2025mtl 2.69% 80.55 88.38 75.26 85.80 88.41 80.45 95.57 81.61 80.75 84.09
LoRAMoE dou2024loramoe 2.19% 80.97 88.51 74.37 86.20 89.45 80.91 95.26 81.29 82.06 84.34
[…]
Table 6: Results (%) of various methods on NLU-MTL. The best results on each dataset are shown in bold, and the second best results are shown in underline. We report the matched accuracy for MNLI, Matthew’s correlation for CoLA, and average correlation for STS-B.
Method #Params CoLA MNLI MRPC QNLI QQP RTE SST-2 Overall
LoRA hu2021lora 2.09% 39.69 86.20 80.88 88.16 87.80 80.14 90.60 79.07
MixLoRA li2024mixlora 3.00% 61.71 89.98 82.84 94.47 90.11 85.56 95.53 85.74
MoSLD zhao2025mosld 1.49% 58.55 90.34 85.29 94.95 90.08 89.53 96.33 86.44
MTL-LoRA yang2025mtl 2.69% 60.65 90.26 87.01 95.26 91.95 90.98 96.10 87.46
HydraLoRA tian2024hydralora 2.72% 64.92 89.97 86.76 95.31 91.62 87.73 96.44 87.54
LoRAMoE dou2024loramoe 2.19% 62.84 91.01 88.97 95.55 91.01 90.61 96.56 88.08
[…]
4.2 Performance in Conflict-Resistance
Tables 5 and 6 compare MoORE with its competitors on CSR-MTL and NLU-MTL, respectively. Using a comparable number of learnable parameters for each dataset, MoORE achieves the best or comparable performance across all tasks and thus obtains the best average performance. These results demonstrate the superiority of MoORE in mitigating task conflict.
The impact of orthogonal adapter. In the experiment on CSR-MTL, we increase the number of Householder reflections in $\bm{H}$ (i.e., $L$) from 0 to $8$ and find that MoORE exhibits consistent improvements in performance. In the experiment on NLU-MTL, however, applying the orthogonal adapter may not improve performance. In our opinion, this phenomenon indicates that the commonsense reasoning tasks in CSR-MTL require more domain knowledge not covered by the pre-trained model. As a result, introducing the orthogonal adapter increases the number of learnable parameters and enhances the model capacity accordingly. In contrast, the text classification tasks in NLU-MTL rely more on the non-specific natural language knowledge captured by the pre-trained model. Therefore, without introducing more learnable parameters, adjusting the singular values of the pre-trained weight matrix is sufficient to achieve encouraging performance.
Conflict-resistance regarding task number and difficulty. To further compare and analyze the conflict-resistance capabilities of different methods, we conduct comparative experiments on CSR-MTL by varying the number and difficulty of the tasks. In particular, for each task in CSR-MTL, we first calculate the average of all the methods’ performance based on the results in Table 5. The average performance of the methods in a task measures the difficulty of the task for the base model — the lower the average performance is, the more difficult the task is. Then, we sort the tasks in ascending order based on their difficulty. Finally, we adapt the base model for the top-$K$ tasks, K=2,,9K=2,...,9, and show the performance of different methods in Figure 1(b). With the increase of task number and difficulty, all the methods suffer performance degradation because $i)$ task conflict becomes severe as the number of tasks increases, and ii)ii) difficult tasks are more likely to have conflicts with other tasks. Notably, MoORE outperforms all other baselines across all settings.
(a) MMLU
(b) IFEval
(c) BBH
(d) GPQA
(e) HumanEval
(f) MBPP
(g) GSM-8K
(h) Overall degradation
Figure 2: The loss of performance in OR-MTL. (a-g) The gray dashed line represents the performance of LLaMA-3.1 8B before adaptation. (h) The overall performance degradation across all tasks.
4.3 Performance in Oblivion-Resistance
To demonstrate the superiority of MoORE in mitigating task oblivion, we compare various methods in the Language Model Evaluation Harness framework eval-harness . In particular, we first adapt the base model on CSR-MTL using different methods. Then, we evaluate the adapted models on OR-MTL, comparing them with the base model. Figure 2 shows that MoORE consistently mitigates task oblivion across all seven tasks of OR-MTL, with an average performance drop of only 1.31% compared to the base model. It significantly outperforms the other adaptation methods. On HumanEval, MoORE achieves performance exceeding the original model, with an improvement of 9.75%. Similarly, LoRAMoE and MTL-LoRA also obtain slight improvements of 4.27% and 0.16%, respectively. This intriguing phenomenon may imply that the datasets in CSR-MTL have some relevance to HumanEval, providing information and capabilities beneficial for solving this task.
Oblivion-resistance regarding task number and difficulty. Furthermore, we conduct experiments to investigate how the ability of oblivion-resistance changes with increased task number and difficulty. The results in Figure 1(c) show that after adapting to the top-3 tasks of CSR-MTL, MoORE and its strongest competitor, LoRAMoE dou2024loramoe , achieve comparable performance on OR-MTL. However, with the increase of task number and difficulty, the performance of LoRAMoE changes slightly while the performance of MoORE increases and becomes superior to that of LoRAMoE. This result demonstrates that MoORE has stronger oblivion-resistance than its competitors.
The impact of task correlation. We investigate the reason for MoORE’s oblivion-resistance empirically. In particular, we consider the samples of several sub-tasks in MMLU and those of six tasks in CSR-MTL. For each task/sub-task, we compute the average weights of MoORE’s experts over its samples, i.e., $\bm{g}_{k}=\frac{1}{|\mathcal{D}_{k}|}\sum_{\bm{x}^{(k)}\in\mathcal{D}_{k}}g(\bm{x}^{(k)})$. Given a sub-task of MMLU and a task of CSR-MTL, denoted as $\mathcal{D}_{k}$ and $\mathcal{D}_{k^{\prime}}$, respectively, we measure their correlation by $\|\bm{g}_{k}-\bm{g}_{k^{\prime}}\|_{2}$. For each sub-task of MMLU, we record the performance degradation of MoORE compared to the base model. Figure 3 shows normalized performance degradation and task correlation. This visualization result indicates that the oblivion-resistance arises from the correlation of tasks — for correlated tasks, the model can learn some common domain knowledge during adaptation and thus avoid catastrophic forgetting.
The generalization to unseen task IDs. The computation of task-level weights requires task IDs as input, but the ID of a new task is unavailable during inference. Therefore, we use the average of the existing task ID embeddings as the embedding of the new task. Specifically, after adapting the base model on CSR-MTL, MoORE only learns the task IDs of the nine tasks in CSR-MTL. During evaluation on OR-MTL, we use the average of the nine task ID embeddings as the task ID embedding of each new task. The results in Figure 1(c) and 2 validate the effectiveness of this approach.
In addition, this experiment also explains the result in Figure 1(c). In particular, the more tasks considered in the adaptation phase, the more likely some tasks correlate with those covered in the pre-training phase. As a result, increasing the number of tasks in the adaptation phase helps enhance the oblivion-resistance of MoORE.
4.4 Inference Efficiency
To compare inference efficiency, we adapt the base model using MoORE and other baselines, and evaluate their inference time on GPQA. For a fair comparison, all experiments are conducted on a single NVIDIA A100 GPU with the same batch size. The results in Figure 1(d) show the average inference time per batch for each method, which indicates that MoORE achieves superior inference efficiency compared to most baselines and is comparable to MixLoRA. Compared to the base model (i.e., the gray dashed line), MoORE moderately increases inference time.
Figure 3: The visualization of normalized performance degradation and task correlation. The “difference” shown in the first row is the normalized performance degradation, i.e., $(\text{Acc}_{\text{Base}}-\text{Acc}_{\text{MoORE}})/100\%$. The following matrix records the normalized task correlation. The element in the $j$-th row and the $i$-th column is $\|\bm{g}_{i}-\bm{g}_{j}\|_{2}/{\max_{k,k^{\prime}}\|\bm{g}_{k}-\bm{g}_{k}^{\prime}\|_{2}}$.
5 Conclusion and Future Work
In this paper, we propose a simple but effective model MoE-ization strategy for multi-task adaptation, leading to a novel MoE architecture, called MoORE. Experiments on various datasets show the effectiveness and superiority of MoORE, demonstrating that imposing orthogonality on experts and maintaining the column space of the pre-trained weight matrix help improve the resistance of the adapted model to task conflict and oblivion.
Limitations and future work. As shown in Figure 1(d), the inference efficiency of MoORE is comparable to that of baselines. However, further reducing its complexity for large-scale applications is challenging. In particular, MoORE applies many more experts than classic MoE models, and the analytic experiments in Appendix C.4 show that it learns dense weights for the experts. In addition, besides imposing orthogonal adapters, we need to find a more effective solution to enhance model capacity in broader scenarios. In theory, the results in Figures 1(b) and 1(c) show that increasing the number of tasks leads to a higher risk of task conflict while helping mitigate task oblivion. How to understand this “trade-off”? How can the limitation in multi-task adaptation be quantified regarding the number of tasks? These are still open problems. Based on the above analysis, we plan to explore efficient implementations of MoORE and study its theoretical properties in the future.
Acknowledgments and Disclosure of Funding
This work was supported by National Natural Science Foundation (92270110), Beijing Natural Science Foundation (L233008), the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China. We also acknowledge the support provided by the fund for building world-class universities (disciplines) of Renmin University of China and by the funds from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, and from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, Renmin University of China.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our claims made in the abstract and introduction accurately reflect the contributions and scope of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of the proposed method in Limitations, such as its inability to fully leverage the potential of orthogonal fine-tuning.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide the complete proof process in both the main text and Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The detailed hyperparameter configurations as well as experimental setups are thoroughly documented in the paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have released the code on an anonymous GitHub repository. The datasets used in the experiments are all publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The detailed hyperparameter configurations as well as experimental setups are thoroughly documented in the paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We did not have sufficient computational resources to run multiple rounds of experiments to obtain error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have provided detailed information about the computational resources used in the experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Every aspect of the paper complies with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have discussed both potential positive societal impacts and negative societal impacts of the work performed in Section D.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the creators or original owners of assets including code, data, and models used in the paper are properly credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the comments in the released code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The proposed method is specifically designed for multi-task adaptation of LLMs, making LLMs an essential component of this paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Hyperparameters and Implementation Details
The detailed hyperparameter setups are presented in Table 7. Both training and testing are conducted on one NVIDIA A100 GPU.
Table 7: Hyperparameter configurations of MoORA on the CSR-MTL and the NLU-MTL.
Hyperparameters MoORA
Cutoff Length 512
Batch Size 8 / 64
Epochs 2 / 5
Learning Rate 3E-04
LR scheduler Warmup-Stable-Decay
Warmup Ratio 5%
Decay Ratio 5%
Optimizer AdamW
Dropout Rate 0.0
Target Modules Q, K, V, O, Up, Down, Gate
$D_{t}$ 128
$D_{s}$ 64
$L$ 8
Appendix B Tasks and Datasets
Detailed information about the CSR-MTL, the NLU-MTL and the OR-MTL is presented in Tables 8, Table 9 and Table 10, respectively. These tables include the sizes of the training and test sets, as well as the task types.
Table 8: The basic information of CSR-MTL.
Task Name #Train #Test Task Type
ARC-Challenge 1,119 1,172 Question Answering
ARC-Easy 2,250 2,380 Question Answering
BoolQ 9,427 3,270 Text Classification
OpenBookQA 4,957 500 Question Answering
PIQA 16,100 1,840 Question Answering
SocialIQA 33,410 1,954 Question Answering
HellaSwag 39,905 10,042 Sentence Completion
WinoGrande 9,248 1,267 Fill in the Blank
CommonsenseQA 9,741 1,140 Question Answering
Table 9: The basic information of NLU-MTL.
Task Name #Train #Validation Task Type
CoLA 8,551 1,043 Text Classification
MNLI 392,702 9,815 -
MRPC 3,668 408 -
QNLI 104,743 5,463 -
QQP 363,846 40,430 -
RTE 67,350 873 -
SST-2 2,491 277 -
Table 10: The basic information of OR-MTL.
Task Name #Test Task Type
MMLU 14,042 Question Answering
IFEval 541 Text Generation
BBH 6,511 Question Answering
GPQA 448 Question Answering
HumanEval 164 Text Generation
MBPP 500 Text Generation
GSM-8K 1,319 Text Generation
Appendix C More Experimental Results
C.1 Performance on other Model Architectures in Conflict-Resistance
To evaluate the cross-model performance of MoORE, we take Qwen-2.5 7B as the backbone model and apply various MoE-based multi-task adaptation methods to it. Experimental results on CSR-MTL are shown in Table 11. We can find that MoORE outperforms baselines consistently under different hyperparameter settings, and the performance of MoORE is robust to the change of $L$. These results verify the robustness of our method and its compatibility with various model architectures.
Table 11: Results (%) of Qwen-2.5 7B adapted by various methods on CSR-MTL. The best results on each dataset are shown in bold, and the second best results are shown in underline.
Method #Params ARC-C ARC-E BoolQ OBQA PIQA SIQA HellaS WinoG CSQA Overall
LoRA [25] 2.12% 87.29 91.88 70.37 90.60 87.92 79.02 92.73 74.98 84.36 84.35
MixLoRA [30] 3.32% 86.86 91.75 71.38 91.40 89.28 80.45 95.06 78.45 85.09 85.53
MoSLD [81] 1.43% 86.69 92.13 71.96 90.40 89.66 80.40 95.17 78.30 85.50 85.58
HydraLoRA [60] 2.83% 88.23 91.54 72.57 91.20 85.80 80.09 95.95 82.24 85.26 85.87
MTL-LoRA [73] 2.80% 87.80 92.09 72.72 92.20 86.07 79.79 95.92 80.66 84.36 85.73
LoRAMoE [14] 2.21% 87.20 91.84 73.61 90.40 90.37 81.27 95.99 80.27 84.60 86.17
[…]
C.2 Numerical Performance in Oblivion-Resistance
Following LLaMA-3.1 8B, we select MMLU [24, 23], IFEval [83], BIG-Bench Hard (BBH) [58], GPQA [51], HumanEval [7], MBPP [4], and GSM-8K [10] as evaluation tasks and adopt the same few-shot settings. More detailed results of the oblivion-resistance experiments are presented in Table 12, corresponding to the results shown in Figures 1(c) and 2.
Table 12: The loss of performance in the original tasks. Each method is represented by two rows of data: the first row indicates the performance on the current task, while the second row shows the difference between the post-fine-tuning score and the pre-fine-tuning score.
Category General Reasoning Code Math
Task MMLU IFEval BBH GPQA HumanEval MBPP GSM-8K Overall
#Shots 5 0 0 0 3 8
Llama-3.1-8B-Instruct 68.15 84.53 46.09 34.15 56.71 59.00 85.52 62.02
LoRA 32.40 29.14 13.65 27.23 0.00 0.00 2.12 14.93
-35.75 -55.39 -32.44 -6.92 -56.71 -59.00 -83.40 -47.09
LoRAMoE 63.74 78.06 34.16 31.47 60.98 57.40 78.54 57.76
-4.41 -6.47 -11.93 -2.68 4.27 -1.60 -6.98 -4.26
MoSLD 61.77 75.06 34.74 28.57 56.10 53.60 70.05 54.27
-6.38 -9.47 -11.35 -5.58 -0.61 -5.40 -15.47 -7.75
MTL-LoRA 61.55 72.30 31.59 31.92 57.32 51.60 66.57 53.26
-6.60 -12.23 -14.50 -2.23 0.61 -7.40 -18.95 -8.76
HydraLoRA 62.34 69.78 30.27 30.36 54.27 54.80 66.41 52.60
-5.81 -14.75 -15.82 -3.79 -2.44 -4.20 -19.11 -9.42
MixLoRA 64.47 77.82 38.21 30.13 54.27 56.20 74.83 56.56
-3.68 -6.71 -7.88 -4.02 -2.44 -2.80 -10.69 -5.46
MoORE 65.32 82.49 40.21 31.92 66.46 58.20 80.36 60.71
-2.83 -2.04 -5.88 -2.23 9.75 -0.80 -5.16 -1.31
C.3 Ablation Study
To analyze the impact of the task-level routing, sample-level routing, and orthogonal adapter quantitatively, we conduct ablation study on the CSR-MTL dataset. In Table 13, we record the overall performance achieved when only one of these three modules is applied. In particular, without the orthogonal adapter, our method degrades to the MoE with fixed experts. When merely applying the orthogonal adapter without any routing, our method is simplified as the single-task adaptation method HRA [75]. The results show that when only one of these three modules is applied, even with a wider network and more trainable parameters, the model’s performance declines significantly across the entire dataset. These results demonstrate that all these three modules contribute to the overall effectiveness of MoORE.
Table 13: Ablation studies of MoORE on CSR-MTL.   indicates the absence of that component.
Method $D_{t}$ $D_{s}$ $L$ #Params Overall
16 0.14% 79.55
Merely Apply 32 0.29% 81.80
Task-level Routing 64 0.58% 83.06
128 1.16% 83.53
16 0.39% 82.30
Merely Apply 32 0.78% 82.91
Sample-level Routing 64 1.57% 83.57
128 3.13% 83.80
2 0.03% 80.76
Merely Apply 4 0.06% 82.15
Orthogonal Adapter 8 0.12% 82.98
16 0.25% 83.90
MoORE 128 64 8 2.84% 85.11
Figure 4: The mean and variance of routing weights obtained in HellaS.
C.4 Analysis of Routing Weights
To investigate the router’s preferences for the experts in MoORE, we analyze the distribution of routing weights across different tasks in the CSR-MTL dataset. Take the $Q$ module from the first attention layer of the adapted model as an example. For each task, we compute and show the mean (red point) and variance (blue stem) of each expert’s weight over all samples of the task in Figures 4 and 6. In each figure, the mean weight of an expert is shown as a red point, and the blue stem associated with the point shows the standard deviation. These visualization results provide us with some insights in the mechanism behind MoORE:
  • Preference on the originally important experts. Compared to other components, the routing weights corresponding to the principal components of $\bm{W}$ (i.e., the experts corresponding to large singular values) are adjusted more significantly. The routing weights corresponding to the significant experts are distributed relatively symmetrically around the zero scale, whereas the routing weights for the minor components are more concentrated above the zero scale. This indicates that the model tends to focus on adjusting the routing weights of the experts that are important in the pre-training phase. In other words, these originally important experts often maintained their significance when adapting for new task.
  • Based on the mean and variance of the routing weights, the nine tasks in CSR-MTL can be grouped into three categories:
    • Large Mean, Large Variance: Tasks such as Hellaswag fall into this group. MoORE performs well on the Hellaswag. A large mean indicates that the model significantly enhances or discards well-learned pretraining knowledge, while a large variance means that the model can capture substantial differences between samples and assign different experts to them with high flexibility.
    • Small Mean, Large Variance: Tasks such as OBQA, PIQA, SIQA, and CSQA belong to this group. This result indicates that for each of these tasks, its samples have significant diversity, and MoORE needs to activate different experts to handle different samples.
    • Small Mean and Variance: Tasks such as ARC-C, ARC-E, BoolQ, and Winogrande (WinoG) are in this group. Except for ARC-E, MoORE performs poorly on these tasks. A small mean indicates minimal adjustment to pretraining knowledge, and a small variance suggests minor differences between samples. For the relatively simple ARC-E task, the model may have already mastered the ability to solve it, requiring little adjustment. However, for the more challenging tasks in this group, the model may struggle to learn the ability to solve them within the column space spanned by the singular vectors.
(a) OBQA
(b) PIQA
(c) SIQA
(d) CSQA
Figure 5: The mean and variance of routing weights obtained in four QA tasks.
(a) ARC-C
(b) ARC-E
(c) BoolQ
(d) WinoG
Figure 6: The mean and variance of routing weights obtained in the remaining four tasks of CSR-MTL.
Appendix D Broader Impacts
MoORE can effectively enhance the performance of LLMs in multi-task adaptation and facilitate the application of LLMs in real-world scenarios. Similar to other multi-task adaptation methods, such as MixLoRA and HydraLoRA, MoORE may result in some potential negative social impacts, such as the risk of abuse for adapting LLMs in illegal activities. However, these issues are not unique to MoORE — other adaptation methods are also susceptible to similar risks. Addressing these issues thoroughly will require advancements in technology, the establishment of relevant legal frameworks, and shifts in societal perspectives.